OpenStudy (anonymous):

What is the mole fraction of K in 55.3 g KBr dissolved in 350 g of water (H2O)? Atomic masses: K= 39.0 g/mol; Br= 79.9 g/mol H=1.0 g/mol O=16.0 g/mol

OpenStudy (anonymous):

Questions: a) What is the solute’s molar mass? b) Convert grams of solute to moles of solute using molar mass as a conversion factor. c) Determine the solvent’s molar mass. d) Convert grams of solvent to moles of solvent using molar mass as a conversion factor. e) What is the mole fraction of K? Answers: a) 39.0 g/mol + 79.9 g/mol = 118.9 g/mol b) 55.3 g / 118.9 g/mol = 0.465 mol c) (1.0*2) + 16.0 g/mol = 18.0 g/mol d) 350 g / 18.0 g/mol = 19.4 mol e) ?

OpenStudy (anonymous):

mole fraction = na / na +nb na number of moles of KBr nb is the number of moles of H2O

OpenStudy (anonymous):

So it would be 118.09 / 118.0 + 18.0 ?

OpenStudy (anonymous):

118.09 / 118.09 + 18.0**

OpenStudy (anonymous):

118.09/(118.09+19.4) no. of moles of H2O are 19.4 which you have found in d part.